[Brahms-l] Shift report 20050226 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 26 2005 - 07:47:19 EST
Shifter leader was ebj
Shifters were 


1.  RHIC was messing around with the beam for most of this shift.  We ended up with 2 hours of beam.  

2.  I took the two zero field runs at 2.3 degrees.

3. I have started one run with te FS at 6 degrees with the meagnets set to B861.

4.  I only noticed one problem with the TOF2 high voltage.  I could not reset this channel.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 26 07:47:42 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 26 2005 - 07:47:51 EST